OpenStudy (anonymous):

Show an equation and a solution for the problem. Crane A can unload a ship in 10 hours and crane B can unload it in 14 hours. How long will it take the two cranes to unload the ship working together? Round to the nearest tenth.

OpenStudy (anonymous):

I think you should take the mean right?

OpenStudy (anonymous):

Is it 12 hours?

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

I'm a bit doubtful, I think the answer should be a little less than 10 because the cranes work together right, so it should take less time.

OpenStudy (lilg132):

crane a takes 1/10 per hour crane b take 1/14 per hour so together it will take 1/10 + 1/14 = 6/35 per hour so 1/(6/35) = 35/6 = 5.8 hours (to the nearest 10th otherwise it would be 5 5/6)

OpenStudy (mertsj):

x/10 + x/14 = 1